### v4

Abstract
Masala-CHAI is the first fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists. It addresses a long-standing challenge in automating netlist generation for analog circuits within circuit design automation. Automating this workflow could accelerate the creation of finetuned LLMs for analog circuit design and verification. We identify key challenges in this automation and evaluate the multi-modal capabilities of state-of-the-art LLMs, particularly GPT-4, to address these issues. We propose a three-step workflow to overcome current limitations: labeling analog circuits, prompt tuning, and netlist verification. This approach aims to create an end-to-end SPICE netlist generator from circuit schematic images, tackling the long-standing hurdle of accurate netlist generation. Our framework demonstrates significant performance improvements, tested on approximately 2,100 schematics of varying complexity. We open-source this solution for community-driven development.
Index Terms:
Analog Design, Automation, LLM
I Introduction
Large Language Models (LLMs) have received significant attention due to their wide-ranging applications, from text summarization to code generation, and have a growing impact across various fields. For hardware design, LLMs have primarily demonstrated potential in the digital domain. This includes tasks such as Verilog code generation [1, 2, 3, 4], assertion generation [5, 6], bug fixing [7, 8], and electronic design automation (EDA) tool scripting [9, 10]. The success of these domain-tailored LLMs relies on access to large and high-quality datasets. For instance, 75K Verilog files from GitHub were used to train the Verigen code generation model [4]. Data scraped from Verilog textbooks was also used to improve the performance of LLMs for Verilog [4]. Textbooks have also proven useful in other domains, for instance, in systems biology [11] and in understanding protein interactions [12].
Building on the success of LLMs in the digital domain, it is natural to explore their application in the analog domain, specifically, in the automated generation of analog circuits from natural language specifications. Analog circuits are described in SPICE (and its many variants), the industry-standard textual representation for simulating analog circuits. SPICE is a low-level description that defines the interconnections between analog components like resistors, capacitors, inductors, and transistors. Open-source datasets for SPICE, unfortunately, are very limited compared to Verilog code, creating a gap in research in this area [13, 14, 15]. However, analog circuit textbooks and research papers are plentiful and contain a trove of analog circuit diagrams, but these are usually in image (or figure) format. These images must then be manually converted to SPICE netlists, which is painstaking and time-consuming.
To address this issue, a recently proposed method, AMSNet [13], introduced a semi-automated way to create a SPICE dataset from analog circuit figures using the capabilities of multi-modal LLMs like GPT-4o that take both image and text inputs and produce both image and text outputs. However, AMSNet still requires manual annotations of nets to obtain accurate SPICE netlists. Even with manual annotations, AMSNet still fails to extract accurate netlists of complex circuits (the reasons for this are described in Section III).
Contributions: Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated method for SPICE netlist generation from analog circuit schematics. We incorporate various techniques, including (i) fine-tuned object detectors to extract passive and active components, (ii) deep Hough transform priors to identify nets, (iii) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (iv) post-extraction verification. We make four contributions:
  1. 1.
    An empirical case study to understand the limitations of state-of-the-art multi-modal LLMs like GPT-4o (shown in AMSNet [13]) in SPICE netlist extraction from schematics.
  2. 2.
    Build Masala-CHAI, the first fully automated framework using LLMs along with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics present in textbooks and papers.
  3. 3.
    Use Masala-CHAI to collect a dataset of $\sim$2,100 SPICE netlists (along with metadata like figure captions) from a sample textbook. We open-source our flow to the community to further research in this area [16].
  4. 4.
    Fine-tune GPT models using the Masala-CHAI dataset to demonstrate the potential use-case of the dataset in automated SPICE netlist generation from English prompts.
II Related Work
LLMs have gained significant attention in chip design [17]. Significant progress has been made in improving Verilog code generation [1, 2, 3, 4], with studies introducing new methodologies that enhance the quality of generated Verilog code. These advances demonstrate how LLMs can streamline and improve digital hardware design workflows. In addition to code generation, prompt engineering has proven effective in chip design [18, 19, 20]. By using LLMs, researchers have conceptualized and designed complex digital hardware architectures efficiently, facilitating faster and more accurate chip development. Beyond code generation, LLMs have found applications in assistant chatbots, script generation, and bug analysis [10]. Similarly, [9] explores LLMs in planning and executing tasks in the Electronic Design Automation (EDA) flow. The use of LLMs in generating assertions and testbenches for verifying the correctness of Integrated Circuit (IC) designs has also seen notable improvements [5, 6, 8, 7].
[21, 22] have explored schematic-to-netlist generation using ML-based approaches. However, these solutions are not reproducible due to the unavailability of any open-source implementation. Recent advancements have extended LLM applications into analog design. Closed-loop Python code generation for analog circuits using LLM agents has shown promising results [14]. LLMs have also demonstrated the ability to generate various circuit topologies from given specifications [23]. Furthermore, a dataset for exploring SPICE netlist generation was recently released in [13]. However, existing approaches require manual modifications of circuit schematics, which significantly limits their scalability for large-scale applications.
III Key Challenges in SPICE Netlist Extraction
Despite advances in multi-modal LLM, as demonstrated by AMSNet [13], current solutions like GPT-4o do not automatically extract accurate SPICE netlists from circuit schematics. Our work begins with an in-depth analysis of the specific failure modes encountered by GPT-4o. Building on these insights, we propose solutions to overcome these limitations, advancing the field toward a fully automated SPICE netlist extraction process that significantly reduces dependence on manual annotations.
III-A Can GPT-4o detect electrical components accurately?
Figure 1: GPT-4o responses when asked to list all components in three sample circuit schematics.
You are an expert analog designer. You will be provided with a schematic, your task is to follow the below instructions carefully: 1. To identify the NMOS and PMOS MOSFET, follow the instructions carefully. For NMOS, the arrow on the source terminal points outwards from the transistor. For PMOS, the arrow on the source terminal points inward towards the transistor. 2. List all the components correctly.
Figure 2: Prompt Tuning guides the GPT to differentiate between NMOS and PMOS.
When generating a SPICE netlist, accurately identifying all components is essential for designers but challenging for language models. To evaluate GPT-4o’s ability to recognize electrical components in circuit schematics, we curated a set of analog schematics featuring resistors, capacitors, inductors, MOSFETs, and various sources from a popular analog textbook [24]. As illustrated in Fig. 1, GPT-4o often misclassifies components, confusing NMOS and PMOS transistors (left and middle schematics) and even omitting key elements such as PMOS transistors and current sources in some cases. However, GPT-4o demonstrates the ability to identify all components accurately in some cases (final schematic), revealing promise despite its inconsistencies.
Solution 1: Electrical Component Detection using Object Detection Networks - Recognizing the limitations of GPT-4o in this task, we propose a dedicated solution using a CNN-based object detection model, using the strong performance of models such as YOLO [25]. By training a CNN model specifically designed for schematic components, we significantly improve the accuracy in detecting and bounding all circuit elements, a method detailed in more detail in Section IV.
Solution 2: Prompt Tuning - We developed a targeted prompt enhancement strategy to refine LLM component differentiation, especially between NMOS and PMOS transistors. By emphasizing structural differences in prompts (see Fig. 2), we achieved notable improvements in accuracy. Our empirical findings support this approach in Section IV.
III-B Can GPT-4o Properly Connect Circuit Components?
Figure 3: SPICE netlist generated by the GPT-4o. For brevity, we only show the part of the netlist that describes transistors.
Once all components are detected, the next step is ensuring that the components are connected correctly so the resulting netlist accurately reflects the schematic. While GPT-4o successfully maps 2-terminal devices, our study uncovered several critical failure modes that compromise the accuracy of netlist generation. These failure modes are illustrated in Fig. 3. Firstly, GPT-4o incorrectly assumes that intersecting nets are connected, even when no connection exists. This error arises partly from inconsistencies in how different schematic notations define connectivity. Secondly, GPT-4o frequently mixes up MOSFET terminals (drain, gate, source). Despite the drain and source’s electrical equivalence, this misidentification disrupts the final netlist’s correctness. Finally, differential input and output voltage pairs pose significant challenges for GPT-4o. For example, in the left schematic (Fig. 3), while components were correctly identified, the source connections for (M1 and M2) were incorrect, and the gate voltage for (M3 and M4) was wrongly assumed to be ‘Vb1’. Similar errors were observed in the right schematic, where differential output was mishandled, and the drain of M5 was incorrectly mapped.
Solution 1: Prompt Tuning - We enhanced the input prompts by explicitly specifying critical design features such as differential pairs and diode-connected topologies. This approach improved GPT-4o’s ability to map terminals to their corresponding nets for simpler designs. However, the solution’s scalability was limited, as it struggled with larger schematics containing more complex net connections.
Solution 2: Automatic Net Annotation - Leveraging our knowledge of component locations, we developed a framework to annotate nets with unique identifiers. By incorporating these annotations into the prompt (see Section IV), GPT-4o’s performance in translating schematics to SPICE netlists improved significantly. This systematic annotation reduced ambiguities and enhanced the LLM’s ability to correctly interpret and connect components, particularly in complex schematics.
IV Methodology
Figure 4: Dataset preparation flow starting from a PDF document consisting of Schematic and related text. Masala-CHAI extracts the relevant information and generates the SPICE netlist for the schematic.
Figure 5: Schematic $\longrightarrow$ Annotated Schematic. Colors denote different component types. Similarly, different nets are clustered.
This section details the methodology developed in this study, addressing the significant challenges outlined in Section III. Specifically, the accurate generation of SPICE netlists remains a critical bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections. To begin the dataset creation flow, as shown in Fig. 4 (➊), it takes a document in PDF format. After that, it is subdivided into 2 parts: (I) Extraction/Summarization of text present in the document comprises of the description of schematic and Key Performance Indicators (KPI) like gain, linearity, bandwidth, noise, gain/phase margin, etc - This will be used as an additional context for fine-tuning LLMs, Section V-D, shown in Fig. 4 (➏ - ➐); (II) Automatic SPICE Netlist Generation - Masala-CHAI: This step starts with extraction of the schematics/image from the document, then goes through the three-step process: 1) Labeling Analog Circuit, 2) Prompt Tuning, and 3) SPICE netlist verification, to finally generate a SPICE netlist, shown in Fig. 4 (➋ - ➎).
IV-A Labeling Analog Circuit
Fig. 4 (➋) step eases the job of the LLM in writing the SPICE netlist from schematics by detecting components and enabling net annotation. As discussed in Section III, SPICE netlist generation is not straightforward, requiring the designer to feed information alongside the schematics. However, this manual process hinders generating large-scale datasets. Thus, we automate labeling all parts of the schematics.
IV-A1 Detection of Circuit Components
We train YoloV8 [25], a state-of-the-art object localization and classification model, specifically for detecting and classifying circuit components with high precision, similar to prior work [21, 22]. Given an input circuit image, YoloV8 localizes each component by regressing its center coordinates, bounding box parameters, and classification label. The model’s backbone, CSPDarknet53 [26], enhanced with a Cross Stage Partial bottleneck [27], effectively integrates high-level semantic information with low-level spatial details, enabling accurate detection even for small-scale components — a critical requirement for high-performance circuit analysis. This process results in schematic diagrams annotated with bounding boxes for all detected components, as illustrated in Fig. 5 (➋). We utilize YoloV8 on the open-sourced dataset [28]. This dataset comprises approximately 4,300 circuit diagram snapshots annotated with bounding boxes across 12 component classes: AC Source, BJT, Battery, Capacitor, DC Source, Diode, Ground, Inductor, MOSFET, Resistor, Current Source, Voltage Source. We resized images to 640x640 and trained for 1000 epochs with a learning rate of 0.01.
IV-A2 Net Detection
We performed net detection in circuit schematics by leveraging pre-trained Deep Hough Transform Line priors [29], a state-of-the-art approach for line detection in images. Unlike conventional methods, this model operates in the Hough domain, parametrizing line segments in polar coordinates to achieve precise line segmentation. We remove components from the schematic to address challenges posed by line segments within components such as capacitors and MOSFETs, leaving only the lines representing nets, as illustrated in Fig. 5 (➌). This ensures a clean input for subsequent net analysis and eliminates potential sources of ambiguity. Following net detection, we propose a simple yet effective heuristic for clustering line segments into nets. Specifically, we group all line segments into a single cluster if their endpoints fall within a radius of 40 pixels, as demonstrated in Fig. 5 (➍). To ensure reliability, we fine-tune the radius parameter and manually verify the consistency of the clustered nets.
IV-B Prompt Tuning
Prompts are crucial in guiding LLMs to generate precise, high-quality outputs. Crafting effective prompts involves meticulous design of their language, structure, and contextual framing, enabling models to respond accurately and meaningfully to user inputs. Through prompt refinement, designers can precisely control model behavior, ensuring the generation of coherent and domain-specific responses. In this study, we systematically explore diverse prompt designs tailored to enhance understanding of schematics. Our approach demonstrates significant effectiveness, as illustrated in Fig. 2 and Fig. 6. Specifically, prompt in Fig. 6 is used at the stage Fig. 4 (➌), facilitating insightful queries to the GPT model.
You are an expert analog designer. You will be provided with a schematic, your task is to follow the below instructions carefully: 1. List all the components which you can observe from the figure. 2. MOSFET are 3 terminal devices with (drain, gate, source). 3. For each component, look at the net number highlighted in red. 4. To identify the source terminal of a MOSFET, choose the net highlighted in red which is nearest to the arrow of the MOSFET. 5. Write a SPICE netlist.
Figure 6: Prompt Tuning guides the GPT to generate SPICE netlist from Schematics.
IV-C SPICE netlist verification
Our case study revealed that some generated netlists (Fig. 4 (➍)) exhibit issues such as floating nets. Although these issues may pass a simulator check, they pose significant risks for large-scale datasets where manual inspection is infeasible. To mitigate this, we developed a Python-based feedback mechanism (Fig. 4 (➎)) that enables LLMs to identify and self-correct such errors automatically, eliminating the need for manual intervention. This solution establishes a robust, closed-loop verification process, which can be further enhanced with additional netlist verification features. We are also releasing this as part of our open-source contribution [16].
V Results
V-A Dataset Creation
Figure 7: Automatic SPICE Netlist generation using Masala-CHAI.
Figure 8: Distribution of (a) Number of Components, (b) Number of Nodes, (c) Number of MOSFETs, and (d) Number of lines of SPICE code, respectively in the dataset.
We curated a comprehensive dataset by selecting schematics from textbooks [24, 30]. Textbooks were chosen over research papers because they offer high-quality schematic images, clear context, and detailed explanations. This makes them more suitable for our purposes, as research papers often present high-level descriptions with advanced schematics that use abstract blocks for complex components. We collected over 2,100 schematic images and applied our proposed Masala-CHAI flow (Fig. 4) to generate the corresponding SPICE netlists. Alongside this, we collected captions and descriptions as detailed in Section IV. Fig. 7 demonstrates how Masala-CHAI translates a schematic into a SPICE netlist. Since our primary goal is to generate structurally correct netlists, each component is marked with default parameters.
Our dataset is carefully characterized to showcase its diversity and complexity, as shown in Fig. 8: (a) displays the variation in the number of components within the schematics, reflecting the complexity of the circuits - a higher number indicates a more complex design; (b) illustrates the distribution of nodes in the schematics, which corresponds to the connectivity between components. Notably, accurately identifying and mapping MOSFET terminals to the correct nets is a significant challenge for LLMs; (c) focuses on the number of MOSFETs present, indicating increased complexity for SPICE generation as the number rises; (d) presents the distribution of the number of lines in the generated SPICE netlists, further emphasizing the varying levels of complexity encompassed in our dataset. By creating this diverse and challenging dataset, we provide a robust benchmark for evaluating and enhancing the capabilities of LLMs in automated SPICE netlist generation.
V-B Evaluation
To evaluate the performance of our proposed method, we leverage graph-matching concepts to compare the structural similarity between SPICE netlists. Unlike traditional approaches, we parse the netlists to extract components and their connections. We identify components by their types (e.g., Resistor, Capacitor, MOSFET) to ensure that our graph representations remain invariant to varying component labels used during netlist generation. This results in a constructed graph “G”, where nodes represent circuit components and edges represent their interconnections, as illustrated Fig. 9. We use Graph Edit Distance (GED) as a quantitative metric to compare the representations of two SPICE netlists. GED ($G_{1}$,$G_{2}$) calculates the sum of the costs associated with transforming $G_{1}$ into $G_{2}$ through a series of edit operations on nodes and edges, such as insertions, deletions, and substitutions. To enhance the interpretability of our results, we normalize the GED to yield a final metric ranging from 0 to 100. This normalization is achieved by calculating the maximum possible GED, obtained by summing the total number of nodes and edges in both graphs and scaling accordingly. This standardized metric allows for a consistent and meaningful comparison across different circuits. Thus, the similarity score between the two netlists is:
$S=\left(1-\frac{\text{GED}(G_{1},G_{2})}{\text{GED}_{\text{max}}}\right)\times 100\%$ (1)
V-C Verifying against AMSNet [13] netlists
We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset, which includes $\sim$800 circuit diagrams. Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Fig. 4). When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier. This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams. Fig. 9 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋). The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions. These findings strengthen Masala-CHAI accuracy and scalability in generating accurate netlists for various circuit diagrams.
Figure 9: Graph representation of SPICE netlists generated by AMSNet [13] and Masala-CHAI. They are identical.
V-D Finetuning
Finetuning entails adapting a pre-trained model to a specific downstream task [4, 3]. This process entails further training the model on a smaller task-specific dataset, enabling it to adjust its parameters to capture new patterns and subtle nuances relevant to the task without training from scratch. The advantages of fine-tuning are: 1) It substantially reduces training time and computational resources by leveraging the foundational knowledge already embedded in the model; 2) It enhances performance on specialized tasks, particularly when data is scarce; and 3) It facilitates the transfer of learned representations across different domains, thereby improving efficiency and effectiveness in model deployment.
GPT: OpenAI [31] released GPT models up to the state-of-the-art GPT-4o model, which is expensive. Fine-tuning [32] reduces cost and latency by replacing the expensive GPT-4o, with a fine-tuned GPT-4o-mini model. The fine-tuned GPT-4o-mini can achieve quality results similar to those of GPT-4o.
Table I reports preliminary results of fine-tuning the GPT’s models on the Masala-CHAI datasets generated. We picked 20 common design problems frequently used as a standalone circuit or sub-part of a bigger design. We prompted the LLM to generate (n = 10) samples of SPICE netlist for each circuit description and manually validated their correctness (the score reported is the number of correct SPICE netlists). We subdivided the 20 benchmarks into 3 different categories: Easy (ID=1-7), Medium (ID=8-14), and Hard (ID=15-20).  Table I provides a comprehensive evaluation of the performance of GPT-3.5-turbo and GPT-4o-mini, both in their ‘base’ and ‘finetune’ states, on a diverse set of circuit design tasks. Finetuning demonstrates significant performance gains across nearly all circuits, particularly for complex designs. For instance, in the (ID=8), GPT-3.5-turbo improves from a score of 4 to 10. At the same time, GPT-4o-mini achieves a perfect score of 10 after fine-tuning, demonstrating the impact of domain-specific training. This pattern is evident across other tasks such as the (ID=5), where fine-tuning boosts GPT-3.5-turbo from 5 to 9 and GPT-4o-mini from 6 to 10.
The results reveal an important trend: simple circuits, such as the (ID=1) and (ID=3), are well-handled even by the base models. More advanced circuits like the (ID=14) and (ID=17) require fine-tuning to achieve meaningful results. Notably, finetuning enables GPT-4o-mini to score 8 on the (ID=10), doubling its base score of 4, illustrating the model’s scalability for complex tasks. GPT-4o-mini consistently outperforms GPT-3.5-turbo in handling advanced designs, achieving superior scores across several tasks.
Table I reveals significant challenges, particularly with specialized circuits like the (ID=19) and (ID=20), where both models exhibit minimal gains even after finetuning.  Fig. 10 showcases representative examples of both successful and failed netlists generated by LLMs, illustrating critical design challenges and the effectiveness of our approach. Notably, in Fig. 10: (1) demonstrates the improper use of an NMOS as a current source rather than forming the intended cascode configuration; (2) highlights a fundamental error where the gate of a diode-connected transistor, expected to connect to the drain, is incorrectly connected to the source; (3) the complexity of a two-stage design further amplifies the difficulties in understanding and replicating the intended topology. These examples highlight the inherent limitations of unoptimized LLMs. Through the ‘Finetune’ GPT-4o-mini model, we successfully resolved these issues, achieving accurate connections and eliminating floating nets. This demonstrates the value of fine-tuning in enabling reliable and context-aware netlist generation.
Figure 10: Examples of passed and failed design cases generated by the fine-tuned model.
TABLE I: Results on 20 designs show the success rate of SPICE netlist generation (out of 10 samples, manually validated). ‘Base’ refers to the baseline model, and ‘Finetune’ indicates its fine-tuned version.
ID Circuit Description GPT-3.5-turbo GPT-4o-mini
Base Finetune Base Finetune
1 Common-source amplifier 8 10 8 10
2 2-stage common source amplifier with resistive load 6 10 8 10
3 Common-drain amplifier 8 10 10 10
4 common-gate amplifier 7 10 10 10
5 Single-Stage RC Low-Pass Filter 5 9 6 10
6 Source Degenerated Amplifier 6 8 6 10
7 Current Mirror 5 9 5 9
8 Common-source amplifier using active load 4 10 7 10
9 Cascode amplifier using NMOS and resistive load 6 10 6 8
10 1-stage differential amplifier 2 7 4 8
11 Diode-connected Amplifier 0 3 3 9
12 Buffer design using MOSFET 0 4 2 8
13 2-input NAND gate 0 5 0 8
14 2-stage amplifier with miller compensation 0 1 0 7
15 SRAM cell with 6 transistors 0 1 0 6
16 2-stage op-amp with differential inputs and single-handled output 0 0 0 4
17 Fully Differential Amplifier with Common-Mode Feedback 0 0 0 3
18 Cross-coupled LC oscillator 0 0 0 3
19 Telescopic cascode operational amplifier 0 0 0 2
20 Bandgap Reference Amplifier 0 0 0 0
VI Conclusion and Future Work
In this paper, we introduced Masala-CHAI, a novel framework to create large-scale datasets for SPICE netlists from documents comprising analog circuit schematics, addressing critical challenges in the design workflow. Through schematic preprocessing and prompt tuning, we developed a fully automated pipeline that leverages LLMs for scalable netlist generation. To the best of our knowledge, Masala-CHAI is the first LLM-based framework of its kind, and our released dataset and flow [16] provide a valuable resource for the community for future advancements.
Future work will focus on enhancing the framework with automated parameter tuning to ensure functionally validated netlists. Additionally, we aim to extend automation to the layout generation stage, addressing its inherent challenges by curating comprehensive datasets and fine-tuning LLMs for this task. These advancements will further streamline the analog design process, bridging the gap from schematic to GDSII and unlocking new efficiencies in IC design workflows.

### v5

@@
 Abstract
-Masala-CHAI is the first fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists.
-It addresses a long-standing challenge in automating netlist generation for analog circuits within circuit design automation.
-Automating this workflow could accelerate the creation of finetuned LLMs for analog circuit design and verification.
-We identify key challenges in this automation and evaluate the multi-modal capabilities of state-of-the-art LLMs, particularly GPT-4, to address these issues.
-We propose a three-step workflow to overcome current limitations: labeling analog circuits, prompt tuning, and netlist verification.
-This approach aims to create an end-to-end SPICE netlist generator from circuit schematic images, tackling the long-standing hurdle of accurate netlist generation.
-Our framework demonstrates significant performance improvements, tested on approximately 2,100 schematics of varying complexity.
-We open-source this solution for community-driven development.
+Masala-CHAI is a fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists.
+It addresses a long-standing challenge in circuit design automation: automating netlist generation for analog circuits.
+Automating this workflow could accelerate the creation of fine-tuned LLMs for analog circuit design and verification.
+In this work, we identify key challenges in automated netlist generation and evaluate multimodal capabilities of state-of-the-art LLMs, particularly GPT-4, in addressing them.
+We propose a three-step workflow to overcome existing limitations: labeling analog circuits, prompt tuning, and netlist verification.
+This approach enables end-to-end SPICE netlist generation from circuit schematic images, tackling the persistent challenge of accurate netlist generation.
+We utilize Masala-CHAI to collect a corpus of 7,500 schematics that span varying complexities in 10 textbooks and benchmark various open source and proprietary LLMs.
+Models fine-tuned on Masala-CHAI when used in LLM-agentic frameworks such as AnalogCoder achieve a notable 46% improvement in Pass@1 scores.
+We open-source our dataset and code for community-driven development.
 Index Terms:
@@
 The success of these domain-tailored LLMs relies on access to large and high-quality datasets.
-For instance, 75K Verilog files from GitHub were used to train the Verigen code generation model [4].
-Data scraped from Verilog textbooks was also used to improve the performance of LLMs for Verilog [4].
-Textbooks have also proven useful in other domains, for instance, in systems biology [11] and in understanding protein interactions [12].
-Building on the success of LLMs in the digital domain, it is natural to explore their application in the analog domain, specifically, in the automated generation of analog circuits from natural language specifications.
-Analog circuits are described in SPICE (and its many variants), the industry-standard textual representation for simulating analog circuits.
-SPICE is a low-level description that defines the interconnections between analog components like resistors, capacitors, inductors, and transistors.
-Open-source datasets for SPICE, unfortunately, are very limited compared to Verilog code, creating a gap in research in this area [13, 14, 15].
+For example, the Verigen code generation model [4] was trained on a dataset of 75,000 Verilog files sourced from GitHub, with additional data extracted from Verilog textbooks to further enhance its performance in Verilog-related tasks.
+Beyond hardware design, textbooks have also been valuable in other domains, such as systems biology [11] and protein interaction studies [12], demonstrating their broader utility in training domain-specific models.
+Figure 1: Pass@1 performance of GPT-4o fine-tuned with Masala CHAI datasets extracted from between 1-10 textbooks, compared to the no fine-tuning baseline (task IDs from Table II).
+Our largest dataset of  7500 captioned SPICE netlists extracted from 10 textbooks provides more than 40% Pass@1 accuracy even on challenging tasks like telescoping cascode amplifier and bandgap reference circuit generation.
+Building on the success of LLMs in the digital domain, it is natural to extend their application in the analog domain, particularly in the automated generation of analog circuits from natural language specifications.
+Analog circuits are described in SPICE and its various variants, which serve as the industry-standard textual representation for analog circuit simulation.
+SPICE provides a low-level description that defines the interconnections between analog components like resistors, capacitors, inductors, and transistors.
+The availability of open-source datasets for SPICE remains significantly limited compared to Verilog, which poses a major challenge for automating the learning of frameworks for the generation of net lists [13, 14, 15, 16, 17].
 However, analog circuit textbooks and research papers are plentiful and contain a trove of analog circuit diagrams, but these are usually in image (or figure) format.
 These images must then be manually converted to SPICE netlists, which is painstaking and time-consuming.
-To address this issue, a recently proposed method, AMSNet [13], introduced a semi-automated way to create a SPICE dataset from analog circuit figures using the capabilities of multi-modal LLMs like GPT-4o that take both image and text inputs and produce both image and text outputs.
-However, AMSNet still requires manual annotations of nets to obtain accurate SPICE netlists.
-Even with manual annotations, AMSNet still fails to extract accurate netlists of complex circuits (the reasons for this are described in Section III).
-Contributions: Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated method for SPICE netlist generation from analog circuit schematics.
-We incorporate various techniques, including (i) fine-tuned object detectors to extract passive and active components, (ii) deep Hough transform priors to identify nets, (iii) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (iv) post-extraction verification.
-We make four contributions:
+Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated framework for SPICE netlist generation from analog circuit schematics.
+We incorporate various techniques, including (i) Schematic extractor from documents, (ii) fine-tuned circuit component detectors to extract passive and active components, (iii) deep hough transform priors to identify nets, (iv) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (v) post-extraction verification.
+We curated the largest open-source analog circuit dataset, comprising 7,500 examples using Masala-CHAI, and fine-tuned various LLMs, including CodeLlama, DeepSeek-34B, and GPT-(3.5,4).
+As shown in  Figure 1, larger datasets significantly enhance performance, improving pass@k results across different numbers of textbooks used for fine-tuning.
+Our contributions are as follows:
   1. 1.
@@
   2. 2.
-    Build Masala-CHAI, the first fully automated framework using LLMs along with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics present in textbooks and papers.
+    Build Masala-CHAI, the first fully automated framework that combines LLMs with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics found in textbooks and research papers.
   3. 3.
-    Use Masala-CHAI to collect a dataset of $\sim$2,100 SPICE netlists (along with metadata like figure captions) from a sample textbook.
-We open-source our flow to the community to further research in this area [16].
+    Leverage Masala-CHAI to compile an open-sourced dataset of $\sim$7,500 SPICE netlists with metadata (figure captions, components, and nodes) from 10 textbooks.
+To the best of our knowledge, this is the largest curated and labeled SPICE dataset available currently.
   4. 4.
-    Fine-tune GPT models using the Masala-CHAI dataset to demonstrate the potential use-case of the dataset in automated SPICE netlist generation from English prompts.
+    Finetune and benchmark four state-of-art LLMs (CodeLLama-70B, DeepSeek-34B, GPT-3.5, GPT-4o) on the Masala-CHAI dataset, showcasing its potential for automated SPICE netlist generation from English prompts.
+  5. 5.
+[…]
 II Related Work
 LLMs have gained significant attention in chip design [18].
-Significant progress has been made in improving Verilog code generation [1, 2, 3, 4], with studies introducing new methodologies that enhance the quality of generated Verilog code.
+Tremendous progress has been made in improving Verilog code generation [1, 2, 3, 4].
 These advances demonstrate how LLMs can streamline and improve digital hardware design workflows.
 In addition to code generation, prompt engineering has proven effective in chip design [19, 20, 21].
-By using LLMs, researchers have conceptualized and designed complex digital hardware architectures efficiently, facilitating faster and more accurate chip development.
-Beyond code generation, LLMs have found applications in assistant chatbots, script generation, and bug analysis [10].
-Similarly, [9] explores LLMs in planning and executing tasks in the Electronic Design Automation (EDA) flow.
+Using LLMs, researchers have efficiently conceptualized and designed complex digital hardware architectures, facilitating faster and more accurate chip development.
+Beyond code generation, LLMs have found applications in assistive chatbots, script generation, and bug analysis [10].
+[9] explores LLMs in planning and executing tasks in the Electronic Design Automation (EDA) flow.
 The use of LLMs in generating assertions and testbenches for verifying the correctness of Integrated Circuit (IC) designs has also seen notable improvements [5, 6, 8, 7].
@@
 However, these solutions are not reproducible due to the unavailability of any open-source implementation.
-Recent advancements have extended LLM applications into analog design.
-Closed-loop Python code generation for analog circuits using LLM agents has shown promising results [14].
-LLMs have also demonstrated the ability to generate various circuit topologies from given specifications [23].
-Furthermore, a dataset for exploring SPICE netlist generation was recently released in [13].
-However, existing approaches require manual modifications of circuit schematics, which significantly limits their scalability for large-scale applications.
+Recent advances extended LLM applications into analog design.
+Closed-loop Python code generation for analog circuits by LLM agents has shown promising results [14].
+LLMs are able to generate circuit topologies from specifications [22].
+A dataset for exploring SPICE netlist generation was recently released in [13], summarized in Table I.
+TABLE I: Comparison with AMSNet [13]. Here, ‘M’ denotes Manual, ‘P’ denotes Partial Automation (with some manual intervention), and ‘A’ denotes Fully Automated.
+Schematic Component Nets Textbooks Dataset
+AMSNet [13] M A P 1 734
+Masala-CHAI A A A 10 7500
 III Key Challenges in SPICE Netlist Extraction
-Despite advances in multi-modal LLM, as demonstrated by AMSNet [13], current solutions like GPT-4o do not automatically extract accurate SPICE netlists from circuit schematics.
-Our work begins with an in-depth analysis of the specific failure modes encountered by GPT-4o.
-Building on these insights, we propose solutions to overcome these limitations, advancing the field toward a fully automated SPICE netlist extraction process that significantly reduces dependence on manual annotations.
+Despite advances in multi-modal LLMs, models like GPT-4o do not automatically extract accurate SPICE netlists from circuit schematics [13].
+Our work begins with an in-depth analysis of GPT-4o’s failure modes.
+Building on these insights, we propose our solutions and introduce a fully automated SPICE netlist extraction framework without manual annotations.
 III-A Can GPT-4o detect electrical components accurately?
@@
 Figure 3: Prompt Tuning guides the GPT to differentiate between NMOS and PMOS.
-When generating a SPICE netlist, accurately identifying all components is essential for designers but challenging for language models.
-To evaluate GPT-4o’s ability to recognize electrical components in circuit schematics, we curated a set of analog schematics featuring resistors, capacitors, inductors, MOSFETs, and various sources from a popular analog textbook [24].
-As illustrated in Fig.
-1, GPT-4o often misclassifies components, confusing NMOS and PMOS transistors (left and middle schematics) and even omitting key elements such as PMOS transistors and current sources in some cases.
-However, GPT-4o demonstrates the ability to identify all components accurately in some cases (final schematic), revealing promise despite its inconsistencies.
-Solution 1: Electrical Component Detection using Object Detection Networks - Recognizing the limitations of GPT-4o in this task, we propose a dedicated solution using a CNN-based object detection model, using the strong performance of models such as YOLO [25].
-By training a CNN model specifically designed for schematic components, we significantly improve the accuracy in detecting and bounding all circuit elements, a method detailed in more detail in Section IV.
+When generating a SPICE netlist, accurately identifying all circuit components is essential for designers and challenging for language models.
+We curated a set of analog schematics featuring resistors, capacitors, inductors, MOSFETs, and various sources from a popular analog textbook [23] and evaluated GPT-4o’s ability to recognize electrical components in circuit schematics.
+As shown in Figure 2, GPT-4o frequently misidentifies components, often confusing NMOS and PMOS transistors (left and middle schematics) or omitting key elements like PMOS transistors and current sources.
+However, in some cases (final schematic), it accurately identifies all components, demonstrating potential despite its inconsistencies.
+Solution 1: Customized Object Detection Network for Analog Circuit Components- Recognizing the limitations of GPT-4o in this task, we propose a dedicated solution using a deep CNN-based object detection model, using the strong performance of models such as YOLO [24].
+By training a CNN model designed explicitly for schematic components, we significantly improve the accuracy in detecting and bounding all circuit elements, a method detailed in more detail in Section IV.
 Solution 2: Prompt Tuning - We developed a targeted prompt enhancement strategy to refine LLM component differentiation, especially between NMOS and PMOS transistors.
-By emphasizing structural differences in prompts (see Fig.
-2), we achieved notable improvements in accuracy.
+By emphasizing structural differences in prompts (see Figure 3), we achieved notable improvements in accuracy.
 Our empirical findings support this approach in Section IV.
-III-B Can GPT-4o Properly Connect Circuit Components?
+III-B Can GPT-4o accurately connect circuit components?
 Figure 4: SPICE netlist generated by the GPT-4o. For brevity, we only show the part of the netlist that describes transistors.
-Once all components are detected, the next step is ensuring that the components are connected correctly so the resulting netlist accurately reflects the schematic.
+After detecting all circuit components, the next step is to ensure their correct connections, so the generated netlist accurately represents the schematic.
 While GPT-4o successfully maps 2-terminal devices, our study uncovered several critical failure modes that compromise the accuracy of netlist generation.
-These failure modes are illustrated in Fig.
-3.
+These failure modes are illustrated in Figure 4.
 Firstly, GPT-4o incorrectly assumes that intersecting nets are connected, even when no connection exists.
-This error arises partly from inconsistencies in how different schematic notations define connectivity.
-Secondly, GPT-4o frequently mixes up MOSFET terminals (drain, gate, source).
-Despite the drain and source’s electrical equivalence, this misidentification disrupts the final netlist’s correctness.
+This error arises partly from inconsistencies in the definition of connectivity across different schematic notations.
+Secondly, GPT-4o often misidentifies MOSFET terminals, confusing the drain, gate, and source.
+Despite the drain and source’s electrical equivalence, this misidentification affects the accuracy of the final netlist.
 Finally, differential input and output voltage pairs pose significant challenges for GPT-4o.
-For example, in the left schematic (Fig.
-3), while components were correctly identified, the source connections for (M1 and M2) were incorrect, and the gate voltage for (M3 and M4) was wrongly assumed to be ‘Vb1’.
-Similar errors were observed in the right schematic, where differential output was mishandled, and the drain of M5 was incorrectly mapped.
-Solution 1: Prompt Tuning - We enhanced the input prompts by explicitly specifying critical design features such as differential pairs and diode-connected topologies.
-This approach improved GPT-4o’s ability to map terminals to their corresponding nets for simpler designs.
-However, the solution’s scalability was limited, as it struggled with larger schematics containing more complex net connections.
-Solution 2: Automatic Net Annotation - Leveraging our knowledge of component locations, we developed a framework to annotate nets with unique identifiers.
-By incorporating these annotations into the prompt (see Section IV), GPT-4o’s performance in translating schematics to SPICE netlists improved significantly.
-This systematic annotation reduced ambiguities and enhanced the LLM’s ability to correctly interpret and connect components, particularly in complex schematics.
+For example, in the left schematic of Figure 4, while components were correctly identified, the source connections for (M1 and M2) were incorrect, and the gate voltage for (M3 and M4) was wrongly assumed to be ‘Vb1’.
+Similar errors appeared in the right schematic, where differential output was mishandled, and M5’s drain was incorrectly mapped.
+Solution: Automatic Net Annotation using Deep Hough Transform - Leveraging our knowledge of component locations, we developed a method to identify and annotating nets using a Deep Hough transform line priors.
+Incorporating these annotations into the prompt (see Section IV), led to improvement in GPT-4o’s performance in translating schematics to SPICE netlists.
+This systematic annotation reduced ambiguities and enhanced the LLM’s ability to interpret and connect components, particularly in complex schematics correctly.
 IV Methodology
@@
 Figure 6: Schematic $\longrightarrow$ Annotated Schematic. Colors denote different component types. Similarly, different nets are clustered.
-This section details the methodology developed in this study, addressing the significant challenges outlined in Section III.
-Specifically, the accurate generation of SPICE netlists remains a critical bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections.
-To begin the dataset creation flow, as shown in Fig.
-4 (➊), it takes a document in PDF format.
-After that, it is subdivided into 2 parts: (I) Extraction/Summarization of text present in the document comprises of the description of schematic and Key Performance Indicators (KPI) like gain, linearity, bandwidth, noise, gain/phase margin, etc - This will be used as an additional context for fine-tuning LLMs, Section V-D, shown in Fig.
-4 (➏ - ➐); (II) Automatic SPICE Netlist Generation - Masala-CHAI: This step starts with extraction of the schematics/image from the document, then goes through the three-step process: 1) Labeling Analog Circuit, 2) Prompt Tuning, and 3) SPICE netlist verification, to finally generate a SPICE netlist, shown in Fig.
-4 (➋ - ➎).
+This section details the methodology used, addressing the challenges outlined in Section III.
+Accurate generation of SPICE netlists remains a bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections.
+Our dataset creation framework, shown in Figure 5, starts by paginating a textbook into individual PDFs (➊), which are then passed through a YOLOv8 [24], a state-of-the-art object localization and classification model, to extract schematics from the page.
+Along with extracted schematics, we also extract and summarize text describing the schematic, for example, from the schematic caption or text on the page referring to the schematic.
+These are used as additional context for fine-tuning LLMs, Section V-B, shown in Figure 5 (➐ - ➒).
+The extracted schematic is still in image format and needs to be reliably converted to a SPICE netlist.
+To this end, Masala-CHAI follows a three-step process: (i) labeling analog circuits, (ii) prompt tuning, and (iii) SPICE netlist verification, starting from schematic extraction ( Figure 5, ➌ - ➏).
+These steps are described next.
 IV-A Labeling Analog Circuit
-Fig.
-4 (➋) step eases the job of the LLM in writing the SPICE netlist from schematics by detecting components and enabling net annotation.
+Figure 5 (➌) simplifies the LLM’s task of generating SPICE netlists by detecting components and facilitating net annotation.
 As discussed in Section III, SPICE netlist generation is not straightforward, requiring the designer to feed information alongside the schematics.
 However, this manual process hinders generating large-scale datasets.
-Thus, we automate labeling all parts of the schematics.
+Thus, we automate the labeling of all parts of the schematics.
 IV-A1 Detection of Circuit Components
-We train YoloV8 [25], a state-of-the-art object localization and classification model, specifically for detecting and classifying circuit components with high precision, similar to prior work [21, 22].
-Given an input circuit image, YoloV8 localizes each component by regressing its center coordinates, bounding box parameters, and classification label.
-The model’s backbone, CSPDarknet53 [26], enhanced with a Cross Stage Partial bottleneck [27], effectively integrates high-level semantic information with low-level spatial details, enabling accurate detection even for small-scale components — a critical requirement for high-performance circuit analysis.
-This process results in schematic diagrams annotated with bounding boxes for all detected components, as illustrated in Fig.
-5 (➋).
-We utilize YoloV8 on the open-sourced dataset [28].
-This dataset comprises approximately 4,300 circuit diagram snapshots annotated with bounding boxes across 12 component classes: AC Source, BJT, Battery, Capacitor, DC Source, Diode, Ground, Inductor, MOSFET, Resistor, Current Source, Voltage Source.
-We resized images to 640x640 and trained for 1000 epochs with a learning rate of 0.01.
+We train a custom YoloV8 model [24] specifically for detecting and classifying circuit components with high precision, similar to prior work [16, 17].
+YoloV8’s Darknet53 backbone [25], enhanced with a “Cross Stage Partial” bottleneck [26], integrates high-level semantic information with low-level spatial details, enabling accurate detection even for small components—a requirement for component detection in complex circuit schematics.
+We fine-tune YOLOv8 on an open-source dataset [27] consisting of  4,300 circuit diagrams with annotated bounding boxes across 12 component classes: AC Source, BJT, Battery, Capacitor, DC Source, Diode, Ground, Inductor, MOSFET, Resistor, Current Source, and Voltage Source.
+Images were resized to 640x640 and trained for 1000 epochs with a learning rate of 0.01.
+The resulting fine-tuned YoloV8 model localizes each component in the schematic by outputting its center coordinates, a bounding box, and classification labels, as illustrated in Figure 6 (➋).
 IV-A2 Net Detection
@@
 Unlike conventional methods, this model operates in the Hough domain, parametrizing line segments in polar coordinates to achieve precise line segmentation.
-We remove components from the schematic to address challenges posed by line segments within components such as capacitors and MOSFETs, leaving only the lines representing nets, as illustrated in Fig.
-5 (➌).
+We remove components from the schematic to address challenges posed by line segments within components such as capacitors and MOSFETs, leaving only the lines representing nets, as illustrated in Figure 6 (➌).
 This ensures a clean input for subsequent net analysis and eliminates potential sources of ambiguity.
 Following net detection, we propose a simple yet effective heuristic for clustering line segments into nets.
-Specifically, we group all line segments into a single cluster if their endpoints fall within a radius of 40 pixels, as demonstrated in Fig.
-5 (➍).
-To ensure reliability, we fine-tune the radius parameter and manually verify the consistency of the clustered nets.
+Specifically, we group all line segments into a single cluster if their endpoints fall within a radius of 40 pixels, as demonstrated in Figure 6 (➍).
+To ensure reliability, we tune the radius parameter and manually verify the consistency of the clustered nets.
 IV-B Prompt Tuning
@@
 In this study, we systematically explore diverse prompt designs tailored to enhance understanding of schematics.
-Our approach demonstrates significant effectiveness, as illustrated in Fig.
-2 and Fig.
-6.
-Specifically, prompt in Fig.
-6 is used at the stage Fig.
-4 (➌), facilitating insightful queries to the GPT model.
+Our approach demonstrates significant effectiveness, as illustrated in Figure 3 and Figure 7.
+Specifically, prompt in Figure 7 is used at the stage Figure 5 (➍), facilitating insightful queries to the GPT model.
 You are an expert analog designer.
 You will be provided with a schematic, your task is to follow the below instructions carefully: 1.
-List all the components which you can observe from the figure.
+List all components which you can observe from the figure.
 2.
-MOSFET are 3 terminal devices with (drain, gate, source).
+MOSFETs are 3 terminal devices with (drain, gate, source).
 3.
-For each component, look at the net number highlighted in red.
+For each component, look at the net number in red.
 4.
-To identify the source terminal of a MOSFET, choose the net highlighted in red which is nearest to the arrow of the MOSFET.
+To identify source terminal of a MOSFET, choose net highlighted in red which is nearest to the arrow of the MOSFET.
 5.
 Write a SPICE netlist.
-Figure 6: Prompt Tuning guides the GPT to generate SPICE netlist from Schematics.
+Figure 7: Prompt GPT to generate SPICE netlist from Schematic.
 IV-C SPICE netlist verification
-Our case study revealed that some generated netlists (Fig.
-4 (➍)) exhibit issues such as floating nets.
-Although these issues may pass a simulator check, they pose significant risks for large-scale datasets where manual inspection is infeasible.
-To mitigate this, we developed a Python-based feedback mechanism (Fig.
-4 (➎)) that enables LLMs to identify and self-correct such errors automatically, eliminating the need for manual intervention.
-This solution establishes a robust, closed-loop verification process, which can be further enhanced with additional netlist verification features.
-We are also releasing this as part of our open-source contribution [16].
+Our case study revealed that some generated netlists (Figure 5 (➎)) still exhibit issues such as floating nets, posing risks large-scale datasets where manual inspection is infeasible.
+To mitigate this, we introduced a final verification step (Figure 5 (➏)) that calls an LLM on the extracted SPICE netlist to automatically identify and correct common errors, for instance, removing floating nets as mentioned above.
+These checks are provided in our open-source code base, and can be further extended to address other sources of error, as required.
 V Results
@@
 Figure 8: Automatic SPICE Netlist generation using Masala-CHAI.
-Figure 8: Distribution of (a) Number of Components, (b) Number of Nodes, (c) Number of MOSFETs, and (d) Number of lines of SPICE code, respectively in the dataset.
-We curated a comprehensive dataset by selecting schematics from textbooks [24, 30].
-Textbooks were chosen over research papers because they offer high-quality schematic images, clear context, and detailed explanations.
-This makes them more suitable for our purposes, as research papers often present high-level descriptions with advanced schematics that use abstract blocks for complex components.
-We collected over 2,100 schematic images and applied our proposed Masala-CHAI flow (Fig.
-4) to generate the corresponding SPICE netlists.
-Alongside this, we collected captions and descriptions as detailed in Section IV.
-Fig.
-7 demonstrates how Masala-CHAI translates a schematic into a SPICE netlist.
-Since our primary goal is to generate structurally correct netlists, each component is marked with default parameters.
-Our dataset is carefully characterized to showcase its diversity and complexity, as shown in Fig.
-8: (a) displays the variation in the number of components within the schematics, reflecting the complexity of the circuits - a higher number indicates a more complex design; (b) illustrates the distribution of nodes in the schematics, which corresponds to the connectivity between components.
-Notably, accurately identifying and mapping MOSFET terminals to the correct nets is a significant challenge for LLMs; (c) focuses on the number of MOSFETs present, indicating increased complexity for SPICE generation as the number rises; (d) presents the distribution of the number of lines in the generated SPICE netlists, further emphasizing the varying levels of complexity encompassed in our dataset.
-By creating this diverse and challenging dataset, we provide a robust benchmark for evaluating and enhancing the capabilities of LLMs in automated SPICE netlist generation.
-V-B Evaluation
-To evaluate the performance of our proposed method, we leverage graph-matching concepts to compare the structural similarity between SPICE netlists.
+Figure 9: Distribution of (a) # Components, (b) # Nodes, (c) # MOSFETs, and (d) #lines of SPICE code in the dataset.
+We curated a comprehensive dataset by automating schematic extraction and SPICE netlist generation from 10 textbooks [29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
+These textbooks were selected for their high-quality schematic images, clear context, and detailed explanations.
+After collecting $\sim$7,500 schematic images, we applied the Masala-CHAI framework (Figure 5) to generate SPICE netlists.
+Alongside, we collected captions and descriptions as detailed in Section IVFigure 8 demonstrates how Masala-CHAI translates a schematic into SPICE.
+Since our goal is to generate structurally correct netlists, each component is marked with default parameters.
+Our dataset is characterized to highlight its diversity and complexity in  Figure 9: (a) shows the variation in the number of components within schematics, where a higher count indicates greater circuit complexity; (b) presents the distribution of nodes, representing component connectivity.
+Accurately identifying and mapping MOSFET terminals to the correct nets is a significant challenge for LLMs; (c) focuses on the number of MOSFETs, where an increase in count corresponds to higher complexity in SPICE netlist generation; Finally, (d) illustrates the distribution of SPICE netlist line counts, emphasizing the circuit complexities.
+This characterization provides insights into the dataset, aiding in the evaluation of automated SPICE generation techniques.
+By curating this diverse and complex dataset, we establish a robust benchmark for evaluating and improving LLM performance in automated SPICE netlist generation.
+For more analysis see Appendix A-A and  A-B.
+TABLE II: Evaluation benchmarks. We refer each design by its ID.
+ID Circuit ID Circuit ID Circuit
+1 Common-source amplifier 2 2-stage common src amplifier w/ res. load 3 Common-drain amplifier
+4 Common-gate amplifier 5 Single-Stage RC Low-Pass Filter 6 Src Degenerated Amplifier
+7 Current Mirror 8 Common-source amplifier using active load 9 Cascode amplifier using NMOS and res. load
+10 1-stage differential amplifier 11 Diode-connected Amplifier 12 Buffer design using MOSFET
+13 2-input NAND gate 14 2-stage op-amp with Miller compensation 15 SRAM cell with 6 transistors
+16 2-stage op-amp with diff. i/p and single-handled o/p 17 Fully Diff. Amplifier w/ Common-Mode Feedback 18 Cross-coupled LC oscillator
+19 Telescopic cascode op amplifier 20 Bandgap Reference Amplifier
+TABLE III: Results on LLMs and Finetuned (‘FT’) variants on MASALA-CHAI dataset. ‘CL’ is CodeLlama. ‘DS’ is DeepSeek.
+ID CL-70B CL-70B+FT DS-34B DS-34B + FT GPT-3.5 GPT-3.5 + FT GPT-4o GPT-4o + FT
+Pass @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5
+1 13.3 57.1 26.7 84.6 26.7 84.6 40 95.8 66.7 100 93.3 100 66.7 100 100 100
+2 0.0 0.0 13.3 57.1 6.7 33.3 26.7 84.6 53.3 99.3 86.7 100 66.7 100 100 100
+3 26.7 84.6 33.3 91.6 13.3 57.1 46.7 98.1 66.7 100 100 100 80 100 100 100
+4 13.3 57.1 33.3 91.6 26.7 84.6 33.3 91.6 60 99.8 80 100 60 99.8 100 100
+5 6.7 33.3 26.7 84.6 6.7 33.3 26.7 84.6 46.7 98.1 86.7 100 60 99.8 100 100
+6 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 60.0 99.8 73.3 100 80 100 100 100
+7 0.0 0.0 6.7 33.3 0.0 0.0 26.7 84.6 46.7 98.1 86.7 100 53.3 99.3 93.3 100
+8 0.0 0.0 0.0 0.0 0.0 0.0 6.7 33.3 40.0 95.8 66.7 100 46.7 98.1 100 100
+9 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 46.7 98.1 73.3 100 53.3 99.3 86.7 100
+10 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 26.7 84.6 86.7 100 46.7 98.1 93.3 100
+11 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 66.7 100 40 95.8 100 100
+12 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 66.7 100 33.3 91.6 86.7 100
+13 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 53.3 99.3 20.0 73.6 80 100
+14 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 40 95.8 0.0 0.0 73.3 100
+15 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 33.3 91.6 0.0 0.0 86.7 100
+16 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 40 95.8 0.0 0.0 53.3 99.3
+17 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 20 73.6 0.0 0.0 66.7 100
+18 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0 0.0 0.0 0.0 53.3 99.3
+19 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 0.0 0.0 66.7 100
+20 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0 0.0 0.0 0.0 40 95.8
+V-B Finetuning
+Fine-tuning LLMs is widely used to adapt models to specific downstream tasks [4, 3].
+The model leverages knowledge acquired from pre-training on large unsupervised language learning tasks, allowing its weights to be adjusted using a smaller, task-specific dataset.
+This enables the model to capture novel patterns and improve task-specific reasoning without training from scratch.
+Fine-tuning LLMs offers several key advantages: (i) Reduced training time and computational costs by leveraging pre-trained foundational knowledge, (ii) Enhanced performance on downstream tasks, particularly beneficial when data is scarce, and (iii) Improved transfer learning across subtasks within a domain, increasing efficiency in model deployment.
+Benchmarking LLMs: We evaluated the capabilities of LLMs in analog circuit design, using both open-source (CodeLlama-70B [39], DeepSeek-V2 [40]) and proprietary (GPT-3.5,4o-mini [41]) models.
+We fine-tuned all baseline LLMs using the corpus collected through the MASALA-CHAI framework.
+Open-source models were fine-tuned on a Nvidia A100 80GB GPU utilizing LoRA [42], while GPT models were fine-tuned using their respective fine-tuning playground [43].
+[…]
+Testing: We selected 20 design problems as shown in  Table II, frequently used as standalone circuits or subcomponents in larger designs, to form our evaluation benchmark.
+LLMs were prompted to generate the corresponding SPICE netlists.
+Discussion: Table III summarizes the Pass@k performance of various LLMs for SPICE netlist generation.
+As anticipated, open-source models (DeepSeek and CodeLlama) underperform relative to the proprietary GPT models, both before and after fine-tuning.
+Nonetheless, fine-tuning on the collected dataset, (ID=1–5) provides a slight increase in the Pass@k for these open-source models, enabling coverage for additional design IDs (ID=6–8).
+Similarly, fine-tuned GPT-3.5 and GPT-4o achieve higher scores (ID=1–13) and non-zero performance for design IDs (ID=14–20).
+Notably, the fine-tuned GPT-4o (“Best Result”) successfully generates SPICE netlists for all design IDs, highlighting the efficacy of the MASALA-CHAI framework (also illustrated in Figure 1).
+Figure 10 showcases representative examples of both successful and failed netlists generated by LLMs, illustrating critical design challenges and the effectiveness of our approach.
+Notably, in Figure 10: (1) demonstrates the improper use of an NMOS as a current source rather than forming the intended cascode configuration; (2) highlights a fundamental error where the gate of a diode-connected transistor, expected to connect to the drain, is incorrectly connected to the source; (3) the complexity of a two-stage design further amplifies the difficulties in understanding and replicating the intended topology.
+These examples show the limits of unoptimized LLMs.
+Through the finetuned GPT-4o model, we resolved the issues such as eliminating floating nets.
+For more analysis see Appendix A-C.
+Figure 10: Pass/fail design cases generated by fine-tuned model.
+V-C Integration into the AnalogCoder [14] Framework
+AnalogCoder is a feedback-driven LLM framework for automated analog circuit design.
+Table IV reports the improvements in Pass@1 and Pass@5 on Tasks 12-20 by using our fine-tuned GPT-4o (Table III) with AnalogCoder instead of the base GPT-4o model (excluding tasks where AnalogCoder achieves 100% success).
+AnalogCoder’s Pass@1 rate on Task 20 improves from 20% to 60% when used with our fine-tuned GPT-4o model.
+The fine-tuned GPT-4o model alone only achieves a Pass@1 of 40%, suggesting that Masala-CHAI fine-tuning and AnalogCoder feedback together provide best performance.
+TABLE IV: Results with AnalogCoder [14]. Many tasks show improvement and for others, results are the same (100%).
+Task ID AnalogCoder AnalogCoder w/ FT GPT-4o
+Pass@1 Pass@5 Pass@1 Pass@5
+12 13.3 57.1 80 100
+14 73.3 100 100 100
+15 13.3 57.1 86.7 100
+16 6.7 33.3 53.3 99.3
+17 0.0 0.0 60 99.8
+19 60 99.8 80 100
+20 20 73.6 60 99.8
+VI Conclusion
+Masala-CHAI is an automated framework for creating large-scale SPICE netlist datasets from analog textbooks, resulting in the largest publicly available dataset of  7500 SPICE netlists with corresponding captions.
+Fine-tuning LLMs on Masala-CHAI demonstrates its use in automated SPICE netlist generation from natural language prompts.
+Our dataset and codebase (https://masala-chai-llm.github.io/) are fully open-sourced and easily extensible on additional textbooks and documentation.
+Appendix A Appendix
+A-A Evaluation
+Figure 11: Graph representation of SPICE netlists generated by AMSNet [13] and Masala-CHAI. They are identical.
+We leverage graph-matching concepts to compare the structural similarity between SPICE netlists.
 Unlike traditional approaches, we parse the netlists to extract components and their connections.
-We identify components by their types (e.g., Resistor, Capacitor, MOSFET) to ensure that our graph representations remain invariant to varying component labels used during netlist generation.
-This results in a constructed graph “G”, where nodes represent circuit components and edges represent their interconnections, as illustrated Fig.
-9.
-We use Graph Edit Distance (GED) as a quantitative metric to compare the representations of two SPICE netlists.
-GED ($G_{1}$,$G_{2}$) calculates the sum of the costs associated with transforming $G_{1}$ into $G_{2}$ through a series of edit operations on nodes and edges, such as insertions, deletions, and substitutions.
-To enhance the interpretability of our results, we normalize the GED to yield a final metric ranging from 0 to 100.
-This normalization is achieved by calculating the maximum possible GED, obtained by summing the total number of nodes and edges in both graphs and scaling accordingly.
-This standardized metric allows for a consistent and meaningful comparison across different circuits.
+We identify components by types (e.g., Resistor, Capacitor, MOSFET) to ensure that our graph representation is invariant to component labels used during netlist generation.
+This yields a graph “G”, where nodes are circuit components and edges are interconnections, as illustrated Figure 11.
+We use Graph Edit Distance (GED) to compare representations of two SPICE netlists.
+GED ($G_{1}$,$G_{2}$) calculates the sum of the costs associated with transforming $G_{1}$ into $G_{2}$ through a series of edit operations on nodes and edges, such as insert, delete, and substitute.
+We normalize the GED to yield a final metric in the range 0 to 100.
+This normalization is achieved by calculating the maximum GED, obtained by summing the number of nodes and edges in both graphs and scaling accordingly.
+This metric allows for a consistent comparison across different circuits.
 Thus, the similarity score between the two netlists is:
 $S=\left(1-\frac{\text{GED}(G_{1},G_{2})}{\text{GED}_{\text{max}}}\right)\times 100\%$ (1)
-V-C Verifying against AMSNet [13] netlists
-We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset, which includes $\sim$800 circuit diagrams.
-Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Fig.
-4).
+A-B Verifying against AMSNet [13] netlists
+We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset.
+Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Figure 5).
 When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier.
 This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams.
-Fig.
-9 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋).
+Figure 11 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋).
 The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions.
 These findings strengthen Masala-CHAI accuracy and scalability in generating accurate netlists for various circuit diagrams.
-Figure 9: Graph representation of SPICE netlists generated by AMSNet [13] and Masala-CHAI. They are identical.
-V-D Finetuning
-Finetuning entails adapting a pre-trained model to a specific downstream task [4, 3].
-This process entails further training the model on a smaller task-specific dataset, enabling it to adjust its parameters to capture new patterns and subtle nuances relevant to the task without training from scratch.
-The advantages of fine-tuning are: 1) It substantially reduces training time and computational resources by leveraging the foundational knowledge already embedded in the model; 2) It enhances performance on specialized tasks, particularly when data is scarce; and 3) It facilitates the transfer of learned representations across different domains, thereby improving efficiency and effectiveness in model deployment.
-GPT: OpenAI [31] released GPT models up to the state-of-the-art GPT-4o model, which is expensive.
-Fine-tuning [32] reduces cost and latency by replacing the expensive GPT-4o, with a fine-tuned GPT-4o-mini model.
-The fine-tuned GPT-4o-mini can achieve quality results similar to those of GPT-4o.
-Table I reports preliminary results of fine-tuning the GPT’s models on the Masala-CHAI datasets generated.
-We picked 20 common design problems frequently used as a standalone circuit or sub-part of a bigger design.
-We prompted the LLM to generate (n = 10) samples of SPICE netlist for each circuit description and manually validated their correctness (the score reported is the number of correct SPICE netlists).
-We subdivided the 20 benchmarks into 3 different categories: Easy (ID=1-7), Medium (ID=8-14), and Hard (ID=15-20).
-Table I provides a comprehensive evaluation of the performance of GPT-3.5-turbo and GPT-4o-mini, both in their ‘base’ and ‘finetune’ states, on a diverse set of circuit design tasks.
-Finetuning demonstrates significant performance gains across nearly all circuits, particularly for complex designs.
-For instance, in the (ID=8), GPT-3.5-turbo improves from a score of 4 to 10.
-At the same time, GPT-4o-mini achieves a perfect score of 10 after fine-tuning, demonstrating the impact of domain-specific training.
-This pattern is evident across other tasks such as the (ID=5), where fine-tuning boosts GPT-3.5-turbo from 5 to 9 and GPT-4o-mini from 6 to 10.
-The results reveal an important trend: simple circuits, such as the (ID=1) and (ID=3), are well-handled even by the base models.
-More advanced circuits like the (ID=14) and (ID=17) require fine-tuning to achieve meaningful results.
-Notably, finetuning enables GPT-4o-mini to score 8 on the (ID=10), doubling its base score of 4, illustrating the model’s scalability for complex tasks.
-GPT-4o-mini consistently outperforms GPT-3.5-turbo in handling advanced designs, achieving superior scores across several tasks.
-Table I reveals significant challenges, particularly with specialized circuits like the (ID=19) and (ID=20), where both models exhibit minimal gains even after finetuning.
-Fig.
-10 showcases representative examples of both successful and failed netlists generated by LLMs, illustrating critical design challenges and the effectiveness of our approach.
-Notably, in Fig.
-10: (1) demonstrates the improper use of an NMOS as a current source rather than forming the intended cascode configuration; (2) highlights a fundamental error where the gate of a diode-connected transistor, expected to connect to the drain, is incorrectly connected to the source; (3) the complexity of a two-stage design further amplifies the difficulties in understanding and replicating the intended topology.
-These examples highlight the inherent limitations of unoptimized LLMs.
-Through the ‘Finetune’ GPT-4o-mini model, we successfully resolved these issues, achieving accurate connections and eliminating floating nets.
-This demonstrates the value of fine-tuning in enabling reliable and context-aware netlist generation.
-Figure 10: Examples of passed and failed design cases generated by the fine-tuned model.
-TABLE I: Results on 20 designs show the success rate of SPICE netlist generation (out of 10 samples, manually validated).
-‘Base’ refers to the baseline model, and ‘Finetune’ indicates its fine-tuned version.
-ID Circuit Description GPT-3.5-turbo GPT-4o-mini
-Base Finetune Base Finetune
-1 Common-source amplifier 8 10 8 10
-2 2-stage common source amplifier with resistive load 6 10 8 10
-3 Common-drain amplifier 8 10 10 10
-4 common-gate amplifier 7 10 10 10
-5 Single-Stage RC Low-Pass Filter 5 9 6 10
-6 Source Degenerated Amplifier 6 8 6 10
-7 Current Mirror 5 9 5 9
-8 Common-source amplifier using active load 4 10 7 10
-9 Cascode amplifier using NMOS and resistive load 6 10 6 8
-10 1-stage differential amplifier 2 7 4 8
-11 Diode-connected Amplifier 0 3 3 9
-12 Buffer design using MOSFET 0 4 2 8
-13 2-input NAND gate 0 5 0 8
-14 2-stage amplifier with miller compensation 0 1 0 7
-15 SRAM cell with 6 transistors 0 1 0 6
-16 2-stage op-amp with differential inputs and single-handled output 0 0 0 4
-17 Fully Differential Amplifier with Common-Mode Feedback 0 0 0 3
-18 Cross-coupled LC oscillator 0 0 0 3
-19 Telescopic cascode operational amplifier 0 0 0 2
-20 Bandgap Reference Amplifier 0 0 0 0
-VI Conclusion and Future Work
-In this paper, we introduced Masala-CHAI, a novel framework to create large-scale datasets for SPICE netlists from documents comprising analog circuit schematics, addressing critical challenges in the design workflow.
-Through schematic preprocessing and prompt tuning, we developed a fully automated pipeline that leverages LLMs for scalable netlist generation.
-To the best of our knowledge, Masala-CHAI is the first LLM-based framework of its kind, and our released dataset and flow [16] provide a valuable resource for the community for future advancements.
-Future work will focus on enhancing the framework with automated parameter tuning to ensure functionally validated netlists.
-Additionally, we aim to extend automation to the layout generation stage, addressing its inherent challenges by curating comprehensive datasets and fine-tuning LLMs for this task.
-These advancements will further streamline the analog design process, bridging the gap from schematic to GDSII and unlocking new efficiencies in IC design workflows.
+A-C Role of high-quality dataset on Finetuning
+Figure 12: Example showing responses from a fine-tuned GPT-4o model with AMSNet and MASALA-CHAI dataset, respectively.
+As a part of our study, we fine-tuned the GPT-4o with the AMSNet [13] dataset as well and compared the quality of response with the MASALA-CHAI dataset.
+Figure 12 shows the responses for a simple “Common Source Amplifier” (ID=1) from the fine-tuned model using AMSNet and MASALA-CHAI dataset, respectively.
+AMSNet lacks a description of the SPICE netlist present in the dataset, which hinders its performance for supervised fine-tuning (SFT), making it not ready to use for LLM-based tasks that require high-quality description alongside the code for unlocking its full potential.
+MASALA-CHAI also offers such a pipeline to generate description alongside the SPICE netlist from a document, as shown Figure 5 (➑).